Fred is 25 years older than his son. When Fred was the age that his son is now, he was twice as old as his son
Answer by ikleyn(52800)

.
Let F be the Fred's age,
    Then his son's present age is (F-25).


When Fred was at his son's age ?  - 25 years ago.

At that time Fred was (F-25) years and his son was  ((F-25)-25 = (F-50) years old.


The condition says

F- 25 = 2*(F-50).  ====>

F = 75.


Answer.  Fred is 75 years old now.  His son is 50 years old.


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.